OpenStudy (anonymous):

The slope of the straight line passing through the points (0,0) and (3,2) is: 2/3 TRUE OR FALSE?

OpenStudy (smileyxl3):

TRUE

OpenStudy (anonymous):

slope(m) = (y2 - y1) / (x2 - x1) m = (0 - 2) / (0 - 3) m = -2/-3 m = 2/3 TRUE

OpenStudy (anonymous):

thank you very much for your explanation!!!

OpenStudy (anonymous):

no problem :)